# WebQuest: A Benchmark for Multimodal QA on Web Page Sequences

### v1

@@ Abstract
The rise of multimodal LLMs and web agents calls for the creation of challenging benchmarks to evaluate neural architectures. Unlike existing benchmarks that focus on multi-step web navigation, we present WebQuest, a multi-page question-answering dataset that requires simultaneous retrieval and reasoning across web interaction sequences grounded in real-world usage. WebQuest includes three question categories: single-screen reasoning, multi-screen reasoning, and questions based on navigation traces. We evaluate some of the leading multimodal models like GPT-4V, Gemini Flash, and Claude 3 on our dataset, revealing a significant gap between single-screen and multi-screen reasoning. Finally, we investigate inference time techniques like Chain-of-Thought prompting to improve model capabilities on multi-screen reasoning.
WebQuest: A Benchmark for Multimodal QA on Web Page Sequences
Maria Wangthanks: First authors. Correspondence: jdchen@google.com, Srinivas Sunkara11footnotemark: 1, Gilles Baechlerthanks: Equal contribution, Jason Lin22footnotemark: 2, Yun Zhu, Fedir Zubach, Lei Shu, Jindong Chen Google DeepMind
1 Introduction
11todo: 1jdchen: last pass22todo: 2jasonjlin: done
@@ 1 Introduction
Figure 1: An example of multi-page question and answering in WebQuest. In this examples, there are 3 screenshots and 2 question-answer pairs. It requires information extraction from each screenshots and reasoning in order to answer both questions.
As the de facto I/O for Human-Computer Interaction, Graphical User-Interfaces (UIs) allow users to acquire information and complete tasks similar to APIs, as the native interface between computer software. A predominant type of data we consume are multimodal web pages rich in semantic knowledge and structure. The advent of smart devices has promised humans the ease of use, motivating research for multimodal, assistive, predictive, automatic and personalized technology for smart UIs. However, designers of such systems do not always consider that the Internet is modeled after our cognitive patterns, such as a limited attention span and working memory. Bounded by single-page interfaces and the ability to traverse back and forward in a sequence, machine learning techniques have been proposed for UI automation but success on multi-step, multi-screen settings are limited, making them unable to fulfill sophisticated tasks for users. Additionally, there are very few datasets focused on information extraction across navigation traces versus task completion.
Recently, Multimodal Large Language Models (MLLMs) like Gemini Team et al. (2023), GPT-4V OpenAI et al. (2024) have been trained on text and image data crawled from the web to demonstrate powerful zero shot capabilities. However, the data collection process of visual-language datasets Schuhmann et al. (2021, 2022)33todo: 3Which citation(s) we have in mind here? reveals a discrepancy with how a human user would interact with the web: navigating multiple pages searching for information using the various UI elements present on the UI. By simply extracting text and images from web pages, these datasets not only remove the functional congruence shared between desktop and mobile form factors, but also fail to capture the actions taken and the implicit retrieval requirement that is common to a information-seeking or goal-oriented workflow, i.e. flight booking or shopping. As a result, these models are effective in general reasoning, but perform poorly when used as end-to-end autonomous agents.
A key challenge in developing UI agents is that they not only have to interpret spatial relations between structures, they also have to ground open ended instructions within an expanding set of semantically correlated pages across diverse applications. Traditional NLP benchmarks like question answering (QA) and information retrieval have been adapted for visual grounding, but they contain either single-page or multi-page questions that contain the answer on one of the pages. We postulate the lack of a temporal dimension might stem from human’s limited working memory and the limitations of existing models to effectively predict such complex interactions. When provided a sequence of pages, local similarity between adjacent pages requires a model to discern relevant information among confounding concepts. It can therefore be limiting to develop and evaluate LLM agents using only the standard QA formulation. To this end, we modify the QA formulation to include multiple pages, which more accurately reflects a user’s ability to compare and reason across pages, i.e. “If I restrict the take-off time to be between 8:00 AM and 8:00 PM, how much more I need to pay for the best choice in business class?” to arrive at an informed decision or conclusion. We do acknowledge that some sophisticated websites already provide filters and comparison functionalities, but they are limited to popular options and do not always expose all the functionality available to the user.
Motivated by this question, we introduce WebQuest, a multi-page screen question-answering dataset, to emulate both navigation and retrieval reasoning settings formulated as a VQA evaluation of multimodal web agents. The dataset and related code will be released on github at: <to-be-updated>. Figure 1 illustrates an example of multi-screen question answering. We seek to close the gap between general QA-based content understanding and agent model research by providing a novel mode of QA setting. Our dataset targets question-answering across web interaction sequences grounded in real world user behavior. WebQuest includes three question categories: single-screen QA, multi-screen QA, and questions based on navigation traces. The first two involve multimodal content from one or more screens, while navigation trace questions are based on sequences of screens navigated by humans to perform routine tasks. This necessitates the endowment of retrieved knowledge from relevant parts of the sequence to aid with complex screen understanding.
Finally, we evaluate state-of-the-art models GPT-4V OpenAI et al. (2024), Gemini Flash Team (2024), and Claude 3 Anthropic (2024) via prompt engineering and Chain-of-thought prompting, revealing a significant gap between single-page and multi-page reasoning. The main contributions of our work are multifold and we believe it will advance the field of UI understanding and automation.
  • WedQuest is the first dataset containing QA over multiple screens and navigation traces requiring information from more than one screen to answer the question.
@@
  • We release 3 subsets of data containing single-screen, multi-screen and trace images allowing more detailed benchmarking of the capabilities of different models
  • For the trace QA split, we also provide step-by-step supervision for the data which can help improve models on navigation with the aim of answering questions.
2 Related Work
44todo: 4jasonjlin: done
@@ 2 Related Work
The problem of answering questions based on the contents of an image has been studied extensively by the research community from natural image VQA Antol et al. (2015), to TextVQA Singh et al. (2019) of text in natural images and DocVQA Mathew et al. (2021) containing documents with text, tabular structures and figures. Focusing more on documents and UIs, datasets like ChartQA Masry et al. (2022) and InfographicVQA Mathew et al. (2022) emphasize more complex numerical and compositional reasoning. For web and UI understanding, QA datasets like WebSRC Chen et al. (2021) focused on desktop screens and ScreenQA Hsiao et al. (2022) focused on mobile app screens contain questions which require extracting relevant information from a single screen. These two datasets are representative of related work. However, our dataset contains questions that necessitate more complex reasoning and also span multiple screens. Similar to our dataset in having multiple images as input, Multi-page DocVQA Tito et al. (2023) requires visual extraction from up to 20 pages to answer questions. However, Multi-page DocVQA requires identifying a single page that contains the answer whereas in our dataset, the answer spans multiple screens. To the best of our understanding, there are no QA benchmarks spanning multiple websites or images in a task-oriented sequence, where a subset of pages are necessary for joint reasoning.
@@
Computer control and workflow automation involve decision-making grounded in visual inputs. Agents have to plan across diverse APIs, interfaces to execute the correct action at the OS and application level. Datasets either captures grounding in a multi-task setting via natural language expressions Liu et al. (2024) or end-to-end Rawles et al. (2024). Previous modeling focused on text-based observations, while pixel-inputs are increasingly explored in recent work like Shaw et al. (2023), Yang et al. (2023), and Zheng et al. (2024). Despite the increase in datasets related to single-page action grounding and multi-step navigation as discussed in the section below, our dataset is the first to focus on question answering in a single and multi page setting.
@@
Recently datasets that focus on performing various tasks in web and mobile scenarios have garnered widespread attention. Earlier efforts introduced simulated web and mobile environments, such as MiniWob++ Liu et al. (2018), MoTIF Burns et al. (2022), Mind2Web Deng et al. (2023) and WebShop Yao et al. (2023). To facilitate autonomous web agents, WebArena Zhou et al. (2024), VisualWebArena Koh et al. (2024) and WebLinx Lù et al. (2024) pair interactive instructions with DOM trees, HTML and pixel-based environments. A majority of these datasets do not have step-by-step supervision for each action taken as part of completing the task. Recently released MMInA Zhang et al. (2024) is the first to collect multi-hop, sequential navigation across websites. While they focus on long-range reasoning with only one website per task category (e.g. shopping, travel), WebQuest has multiple websites for each task which is closer to real world workflows.
3 WebQuest Benchmark
@@
Figure 2: An example of Single Screen QA, where the task is to count how many cloudy days before a rainy one, and the weather conditions are depicted by pictograms.
Question: How much do all the gym accessories cost in total?
Answer: \$127.77
Figure 3: An example of Multi Screen QA, about prices of selected gym accessories, each from a different online shop. Notice the layouts and font sizes vary among them.
WebQuest is a multimodal benchmark for question answering in a web navigation setting. It is targeted to replicate the scenarios where users navigate to websites to retrieve information across these different websites. To better study the capabilities and failure cases of various modeling approaches, we split these scenarios into three categories:
  1. 1.
    Single screen QA: Questions about the web page the user is currently on. (see Figure 2 for an example);
  2. 2.
@@
    Multi screen QA: Questions across multiple web pages simultaneously displayed to the user in their browser, where the web pages are shortlisted such that all of them are required to answer the question. (see Figure 1 and Figure 3 for examples);
  3. 3.
@@
    Trace QA: Questions about a user’s browsing session containing the entire sequence of screens viewed by the user. (see Figure 13 in the Appendix for an example).
To build a benchmark that represents these three scenarios, we render and capture screenshots of single pages, sets of 3 to 5 relevant pages, and browsing sessions with on average 16 screens, respectively. The questions in this benchmark all involve logical reasoning, requiring models to analyze screen images, extract relevant information from each page, and finally combine the information across pages to answer the question. In this section, we go into details of each of the categories of WebQuest benchmark and discuss the data collection process and then we present detailed statistics of our dataset.
3.1 Single screen Question Answering
The single screen QA benchmark consists of 542 web screenshots with a question-answer pair associated with each screen. To collect the screenshots and associated QA pairs, we provide the raters a list of most popular websites. Then we ask the raters to browse these popular websites, consider the kind of questions a user navigating this website might be interested in, and take a screenshot and generate the relevant QA pair for each such page. The QA pair is required to involve reasoning, including arithmetic (counting, addition, subtraction) and comparison questions. Additionally, we ask the raters to ensure that the questions are diverse and reflect real world scenarios. To address safety and privacy concerns, the raters are asked to only navigate to safe web pages without inappropriate content. They are also instructed to not sign in to any of the websites. Lastly, we do a round of screenshot validations to filter out unsafe screens and ones containing PII. In Figure 4, we present the distribution of different website categories.
Figure 4: Distribution of website categories of Single Screen QA examples.
3.2 Multi screen Question Answering
For multi screen question answering benchmark, we provide 171 QA pairs each containing multiple screens. In the multi screen setting, we aim to capture the case where a user has navigated to several relevant pages simultaneously, most commonly in different tabs, then want to get answers to questions about these different web pages, such as calculating prices of products across websites and comparing different product attributes. To ensure the screens in the same example are relevant, we construct a list of popular websites and group them into different categories, such as shopping, flights, etc. We provide these websites lists to human raters and they are required to navigate to 3 to 5 websites listed in the same category. For example, these sets can be menu pages of several restaurants or flight search pages across different booking websites. After navigating to the web pages, the raters generate the associated question-answer pair with the example. Similar to the single screen questions, the questions are focused on logical reasoning, such as comparisons, counting, filtering, arithmetic, etc. In Figure 5, we present the distribution of number of screens per example. 55todo: 5[Non Urgent]To Maria, for each graph you created in this section, please add a link to the colab or source for it, for future reference.
Figure 5: Distribution of screen lengths of Multi Screen QA examples.
3.3 Trace Question Answering
@@
The trace question answering benchmark consists of 292 examples. The screens in this category represents a sequence of web pages the user navigates to while obtaining the answer to a question. In all cases, only a subset of the screens are relevant to answering the question. As the number of screens increases in the navigation trace, to answer questions based on the screens seen, a representation of the relevant information encountered needs to be maintained. The main difference between this category and the Multi screen QA category is that, in Multi Screen QA each QA pair contains only the screens with information relevant to answer the question, whereas in Trace QA each QA pair also contains the different screens encountered in the same browsing session. We consider Multi screen QA as a simplified version of the Trace QA task which enables us to measure the capabilities of different modeling approaches on answering questions across multiple screens.
For the data collection, similar to the multi-screen screen QA data collection, we provide top websites and their associated categories to the human raters. The raters browse through websites and record screen traces. On average, each trace contains 16 screens. We display the distribution of the number of screens in Figure 6, and the distribution in different categories in Figure 7. Please see Table 4 in the Appendix for exact counts of the websites for the different questions. After constructing the screen navigation history, the rater reviews the screens and generates QA pairs. The raters are instructed to create questions involving multiple screens in the screen sequence and requiring logical or arithmetic reasoning.
Figure 6: Distribution of the length of traces for the Trace QA benchmark. There are a few very long traces with more than 50 screens, while a majority of the traces contain fewer than 20 screens.
@@
Figure 7: Distribution of categories of Trace QA examples. We observe that news, transportation and shopping categories contain the most number of traces.
4 Experiments
inline,color=redinline,color=redtodo: inline,color=redCan you check this?
4.1 Metrics

### v2

@@ Abstract.
The rise of powerful multimodal LLMs has enhanced the viability of building web agents which can, with increasing levels of autonomy, assist users to retrieve information and complete tasks on various human-computer interfaces. It is hence necessary to build challenging benchmarks that span a wide-variety of use cases reflecting real-world usage. In this work, we present WebQuest, a multi-page question-answering dataset that requires reasoning across multiple related web pages. In contrast to existing UI benchmarks that focus on multi-step web navigation and task completion, our dataset evaluates information extraction, multimodal retrieval and composition of information from many web pages. WebQuest includes three question categories: single-screen QA, multi-screen QA, and QA based on navigation traces. We evaluate leading proprietary multimodal models like GPT-4V, Gemini Flash, Claude 3, and open source models like InstructBLIP, PaliGemma on our dataset, revealing a significant gap between single-screen and multi-screen reasoning. Finally, we investigate inference time techniques like Chain-of-Thought prompting to improve model capabilities on multi-screen reasoning.
Datasets,evaluation,WebUIs,Question Answering,Large Language Models
copyright: noneccs: Human-centered computing Human computer interaction (HCI)
1. Introduction
As the de facto I/O for Human-Computer Interaction, Graphical User-Interfaces (UIs) act as a crucial bridge between computers and humans, allowing users to acquire information, complete tasks, engage in social interactions and perform various other daily activities. A predominant type of data we consume are multimodal web pages and apps rich in semantic knowledge and structure. The ubiquity and utility of UIs has motivated research for multimodal, assistive technology to simplify various user journeys. At the same time, recent multimodal LLMs (Team, 2024; OpenAI et al., 2024; Anthropic, 2024) with strong capabilities across a variety of tasks have sparked interest in building end-to-end agents adept at assisting users with various complicated tasks.
One of the challenges in developing UI models is to interpret spatial and semantic relations among the various contents in sequences of screens. Traditional computer vision tasks like summarization, image captioning, visual question answering have been extended to document images (Mathew et al., 2021), infographics (Masry et al., 2022; Mathew et al., 2022) and UIs (Chen et al., 2021; Hsiao et al., 2022; Liu et al., 2024; Wang et al., 2021; Li et al., 2020). However, all of these tasks consist of single-image or multi-image setups that contain the relevant information on at most one of the images. Current SoTA multimodal LLMs achieve impressive performance on these tasks, averaging >85% zero-shot performance across UI and document related datasets like ChartQA (Masry et al., 2022), DocVQA (Mathew et al., 2021), InfographicVQA (Mathew et al., 2022), ScreenQA (Hsiao et al., 2022) and WebSRC (Chen et al., 2021). Motivated by the promise of multimodal LLMs to tackle various UI based single-page tasks, the research community has created a variety of multi page benchmarks focusing on UI Automation (Deng et al., 2023; Koh et al., 2024), completing multi-step tasks on UIs based on user goals. A significant majority of these benchmarks require the use of a single website to complete the given task. In those containing multi-website tasks, typically each example contains one website per category (e.g., shopping, restaurants, movies, education etc) but do not include multiple websites with similar functionality.
Real world web usage necessitates combining and comparing information from multiple webpages. When provided with a set of semantically related webpages, answering challenging questions requires a model to discern relevant information among confounding concepts presented with varied layouts and spatial structures. It can therefore be limiting to develop and evaluate web agents using only the standard QA formulation found in various UI automation datasets. Motivated to address this gap, we introduce WebQuest, a multi-page screen question-answering dataset, to assess information retrieval, reasoning and navigation capabilities of multimodal models.
We reformulate the QA task used in existing UI datasets in two ways. Firstly, we include questions requiring aggregation/reasoning across information extracted from multiple parts of the screen. Secondly, our dataset includes questions spanning multiple webpages and websites, which, we believe, more accurately reflects a user’s ability and need to compare and reason across pages. In figure 1, we provide an example from the dataset, illustrating the same product presented in different websites. In this case, the ability to reason over information extracted from multiple webpages is essential to compare the different offerings and make an informed choice. We acknowledge that some sophisticated websites or aggregator websites provide filters and comparison functionalities, but they are limited to popular options and do not always expose all the functionality available to the user.
To allow for a detailed benchmarking of various models, WebQuest includes three question categories of increasing difficulty: single-screen QA, multi-screen QA, and QA on navigation traces. The first two involve multimodal content from one or more screens, while trace QA is based on sequences of screens navigated by humans to perform routine tasks. Figure 1 illustrates an example of multi-screen question answering. Finally, we evaluate various state-of-the-art proprietary models GPT-4V (OpenAI et al., 2024), Gemini Flash (Team, 2024), Claude 3 (Anthropic, 2024), and open source models like InstructBLIP (Dai et al., 2023), PaLIGemma (Beyer et al., 2024) via prompt engineering and Chain-of-thought prompting, revealing a significant gap between single-page and multi-page reasoning.
@@ 1. Introduction
Figure 1. An example of multi-page question and answering in WebQuest. In this example, there are 4 screenshots and 2 question-answer pairs. Answering both questions needs reasoning over information extracted from the different screenshots.
The main contributions of our work are multifold. We believe it will advance the field of UI understanding and automation.
  • WebQuest is the first dataset containing QA on UIs requiring reasoning across various parts of a single UI or multiple UIs.
  • WebQuest is the first dataset with QA over long interaction sequences focused on information extraction and complex reasoning across the sequences.
@@
  • WebQuest includes 3 subsets of data containing single-screen, multi-screen and navigation traces, enabling detailed benchmarking of the capabilities of different models.
2. Related Work
@@
The problem of answering questions based on the contents of an image has been studied extensively by the research community from VQA (Antol et al., 2015) on natural images, to TextVQA (Singh et al., 2019) of text in natural images and DocVQA (Mathew et al., 2021) containing documents with text, tabular structures and figures. Focusing more on documents and UIs, datasets like ChartQA (Masry et al., 2022) and InfographicVQA (Mathew et al., 2022) emphasize complex numerical and compositional reasoning. For web and UI understanding, QA datasets like WebSRC (Chen et al., 2021) focused on desktop screens and ScreenQA (Hsiao et al., 2022) focused on mobile app screens; both contain questions which require extracting relevant information from a single screen. While related, our dataset contains questions that necessitate more complex reasoning and also span multiple screens. Similar to our dataset in having multiple images as input, Multi-page DocVQA (Tito et al., 2023) requires visual extraction from up to 20 pages to answer questions. However, Multi-page DocVQA requires only a single page that contains the answer to be identified whereas in WebQuest, the answer spans multiple screens. To the best of our understanding, there are no QA benchmarks spanning multiple websites or images in a task-oriented sequence, where a subset of pages are necessary for joint reasoning. Other datasets like TaT-DQA (Zhu et al., 2022) and DUDE (Landeghem et al., 2023) contain questions over multiple pages, but are based on documents with tables and not rich UIs.
@@
Computer control and workflow automation involve decision-making grounded in visual inputs. Agents have to plan across diverse APIs, interfaces to execute the correct action at the OS and application level. Datasets either captures grounding in a multi-task setting via natural language expressions (Liu et al., 2024) or end-to-end (Rawles et al., 2024). Previous modeling focused on text-based abstractions, while pixel-inputs are increasingly explored in recent work like (Shaw et al., 2023), (Yang et al., 2023), and (Zheng et al., 2024). Despite the increase in datasets related to single-page action grounding and multi-step navigation, our dataset is the first to focus on question answering in a single and multi page setting.
@@
Recently datasets that focus on performing various tasks in web and mobile scenarios have received widespread attention. Earlier efforts introduced simulated web and mobile environments, such as MiniWob++ (Liu et al., 2018), MoTIF (Burns et al., 2022), Mind2Web (Deng et al., 2023) and WebShop (Yao et al., 2023). To facilitate autonomous web agents, WebArena (Zhou et al., 2024), VisualWebArena (Koh et al., 2024) and WebLinx (Lù et al., 2024) pair interactive instructions with DOM trees, HTML and pixel-based environments. A majority of these datasets do not have step-by-step supervision for each action taken as part of completing the task. Recently released MMInA (Zhang et al., 2024) is the first to collect multi-hop, sequential navigation across websites. While they focus on long-range reasoning with only one website per task category (e.g. shopping, travel), WebQuest has multiple websites for each task which is closer to real world workflows.
3. WebQuest Benchmark
@@
Figure 2. An example of Single Screen QA, where the task is to count how many cloudy days before a rainy one, and the weather conditions are depicted by pictograms.
Figure 3. Distribution of website categories of Single Screen QA examples.
Question: For nonstop flight from New York City to Hong Kong, how much more expensive is business class than economy?
Answer: \$3054
Figure 4. An example of Trace QA on airfare differences across cabin classes. A browsing session is provided with all screen sequences, but not all of them are required to answer the question.
Figure 5. This figure demonstrates the relationships among Single Screen QA, Multi Screen QA, and Trace QA. Single Screen QA focuses on a single page, Multi Screen QA focuses on multiple pages within a browsing session, and Trace QA focuses on the entire browsing session.
WebQuest is a multimodal benchmark focused on question answering based on the contents of web UIs. It is targeted to capture the scenarios where users navigate to different websites to retrieve information. In contrast with many question answering datasets which are based on information contained in a single image/UI, WebQuest enhances the scope of the QA task by requiring the combination of information from multiple distinct parts of UIs. Furthermore, WebQuest also contains questions spanning multiple related webpages and websites. To better understand the capabilities and characterize the performance of various modeling approaches, we split the dataset into three categories of increasing difficulty:
  1. (1)
    Single screen QA: Questions based on the contents of a single screen (see Figure 3 for an example);
  2. (2)
@@
    Multi screen QA: Questions spanning multiple semantically related web pages. For this task, the web pages are filtered such that all of them are required to answer the question of interest (see Figure 1 for an example);
  3. (3)
@@
    Trace QA: Questions about the entire sequence of screens viewed by the user as they navigate the web with a particular goal (see Figure 4 for an example).
Please refer to figure 5 for an overview of the relation between the different categories. In the remainder of this section, we describe the motivation, details of the different categories in WebQuest and the data collection process. To arrive at the source websites, we initially curate a list of categories based on their functionality and then compile a list of popular websites in each category.
3.1. Single screen QA
This category contains question answer pairs where each question is based on the contents of a single web UI. Compared to other UI based QA datasets (Chen et al., 2021; Hsiao et al., 2022; Liu et al., 2024), answering each question requires extracting information from different parts of the UI and then combining them using mathematical or logical operations. Some of the common cases covered in this splits include counting the number of entities which meet a criteria (e.g., number of items below a certain price), aggregating information etc. This category consists of 542 examples. To collect this dataset, raters start with the list of popular websites provided to them and are asked to follow the instructions:
  1. (1)
    Explore the website and think about what the website is used for and how users typically interact with it.
  2. (2)
    Consider the kind of questions users might be interested in when using the website. Focus on questions using information from multiple parts of the screen and needing arithmetic or logical reasoning.
  3. (3)
    Take a screenshot and generate a relevant QA pair.
During the data collection process, the authors of this work frequently reviewed the generated examples to ensure that the question-answer pairs are of high quality, diverse and reflect real world scenarios. Additionally, two rounds of data validation were performed to remove screenshots that contain PII information. Figure 3 shows the distribution of different website categories in the Single screen QA dataset. We can observe that the Single screen QA dataset contains websites across many categories with shopping and restaurants being the most common ones. The detailed statistics are listed in Appendix section B.
Figure 6. Distribution of categories of Multi Screen QA examples. The websites span over 5 categories with a higher proportion in Shopping and Restaurants.
Figure 7. Distribution of number of screens for Multi Screen QA examples.
3.2. Multi screen QA
In the Multi Screen QA task, we aim to capture instances where users are browsing the web with the aim of gathering information focused on a particular task, e.g., shopping for a sound bar as depicted in fig 3.2. This could involve tasks like comparing product prices/attributes across websites, evaluating options for travel, filtering feasible options for movies etc. The overall goal is similar to the Trace QA described in sec 3.3, however for this case, the screens containing relevant information for answering the questions are already filtered. This occurs in cases where a user has explored several relevant webpages, opened them in different tabs or taken screenshots, and then seeks answers pertaining to these pages. It is important to note that the different screens in each QA pair are semantically related and can come from webpages in the same website or alternative websites offering similar functionality.
This category contains 307 QA pairs each encompassing multiple web screens. The overall guidance provided to raters is similar to the one in section 3.1. One major difference is that now the raters are instructed to capture multiple screenshots from different webpages. To ensure that related websites are explored for each QA pair, raters are provided a curated list of popular websites grouped into categories. This curated list serves as a guide for selecting coherent screens for each QA pair. Each rater is first assigned a category for each question, and are asked to follow the instructions:
  1. (1)
    Choose 2-5 websites from the assigned category.
  2. (2)
    Imagine a user researching a product, service, or location online, comparing information across webpages and websites.
  3. (3)
    Take 3-5 screenshots from websites related to the chosen use case.
  4. (4)
    Formulate a question a user might have based on the combined information in the screenshots.
  5. (5)
    Ensure that the question requires logical or arithmetic operations and needs information from all screenshots to answer.
  6. (6)
    Provide a short answer to the question.
The exact instructions given to the raters are specified in the appendix.
The authors of this work periodically review the data quality and provide feedback to the raters. Furthermore, 2 rounds of data validation were performed to prevent inappropriate and unsafe screens. Figures 7,  7 contain the distribution of different website categories and number of screens per example. Detailed statistics for the categories are presented in section B of the Appendix.
3.3. Trace QA
@@
In the Trace QA category, similar to the Multi Screen QA category described in section 3.2, we aim to capture instances where users are browsing the web with the aim of gathering information focused on a particular task. In the rest of this section, we use the term browsing session to represent the different webpages visited as part of completing this task. All of the webpages encountered during the browsing session are associated with the corresponding QA pair. However, do note that only a subset of the screens are relevant to answering the question. The main difference between this category and the Multi screen QA category is that, in Multi Screen QA each QA pair is associated with only the screens containing information relevant to answer the question, whereas in Trace QA each QA pair contains the different webpages in the browsing setting. We consider Multi screen QA as a simplified version of the Trace QA task which enables us to measure the capabilities of different modeling approaches on answering questions across multiple screens.
For the data collection, similar to the multi-screen screen QA data collection, we provide top websites and their associated categories to the human raters. The raters browse through websites and record screen traces. Below are the instructions provided to the raters: For the Trace QA split, similar to the Multi Screen QA split, the raters are asked to focus on specific website category. The raters install custom Chrome Extension which enables them to record their interactions on the screen. The chrome extension is meant to be turned on when the rater is ready to collect data. They can pause, review or delete any of the data captured. Each rater is first assigned a category for each question, and are asked to follow the instructions:
  1. (1)
    Imagine a user researching a product, service, or location online, comparing information across websites and pages. Consider the kind of questions the user might be interested to know.
  2. (2)
    Go to the first web page and start the trace recording.
  3. (3)
    Navigate through the different webpages as part of the trace. Once the navigation is complete, click "stop" on the recording and review the screenshots captured.
  4. (4)
    Generate the relevant question.
  5. (5)
    Ensure that the question focuses on logical reasoning and arithmetic.
  6. (6)
    Ensure that answering the question needs to use information from about 3-5 screenshots.
We display the distribution of the number of screens in Figure 9, and the distribution in different categories in Figure 9. On average, each trace contains 16 screens. Please see Table 5 in the Appendix for exact counts of the websites for the different questions. The data is validated to ensure no PII is present and the data is manually reviewed for correctness.
@@
Figure 8. Distribution of categories of Trace QA examples. News, transportation and shopping categories contain the most number of traces.
Figure 9. Distribution of the length of traces for the Trace QA benchmark. There are a few very long traces with more than 50 screens, while a majority of the traces contain fewer than 20 screens.
3.4. Rater Details
For the Single Screen QA and Multi Screen QA tasks, we used 10 raters with ages between 25 and 35 based in India. The raters were employed as contractors to help with a variety of data collection tasks apart from this task.
4. Experiments
4.1. Metrics